# Inference of time delay in stochastic systems

### v1

@@ VI Perturbation and deep learning-based approach – numerical results
In the following, we discuss results of the perturbation-based approach to infer time delay using neural networks for the trajectory data obtained from Brownian dynamics simulations.
We will first focus on inference by means of a convolutional neural network trained on the linear time-delayed feedback force [see Eq. (2)]. With this network, we infer time delay for trajectory data obtained from numerical simulations of the linear feedback force [see Eq. (2)] as a proof of concept and subsequently the $\tanh$ feedback force [see Eq. (3)] to demonstrate that the network is able to generalize. In principle, the perturbation-based approach to infer time delay using convolutional neural networks trained on linear feedback forces may also be applicable to other systems that exhibit a stationary or non-equilibrium steady state such as Gaussian time delayed feedback [Eq. (4)]; however, an extensive analysis is beyond the scope of this study.
Finally, we will discuss the inference results obtained from a convolutional neural network trained on trajectories based on Eqs. (12)– (14). We use the trained network to classify trajectory data obtained from numerical simulations of the same equations of motion. In this way, we demonstrate that it is possible to train such a neural network on several functional forms of the (nonlinear) delayed and instantaneous forces simultaneously while maintaining similar prediction accuracy as before. In practical applications, this allows for inference, despite not knowing the exact dynamics beforehand and also without the assumption of a stationary or non-equilibrium steady state.
@@ VI.1 Network trained on linear model applied to linear force data
For our numerical study we first consider the attractive linear feedback force system based on Eqs. (1) and  (2) with $k_{a}=k_{b}>0$ and without additional instantaneous forces. Here we use a neural network that was trained on the same model to infer time delay. In Fig. 10a we show inference results represented by the probability distribution $P(\tau)$ (see Sec. V.2). We averaged over the classification probabilities for different values of the parameters $k_{a}$ and $k_{b}$ at each data point as well as ten noise realizations. Each of the color-coded curves corresponds to one of the exit nodes and represents the probability of a time series to be based on the delay time value $\tau_{\mathrm{node}}$. Markers correspond to the underlying delay times of test data, i.e. the time series we obtain from the perturbation approach and feed into the neural network (after applying the preparation procedure discussed in Appendix A.4). We classify time series based on delay times of $\tau=0\tau_{B}$ (passive system) up to $\tau=0.2\tau_{B}$.
@@
The highest classification probability is observed if a time series is based on a delay time close to one of the $\tau_{\mathrm{node}}$ the network was explicitly trained on. In those cases, we find $P(\tau)\approx 1$, that is, the network is almost certain. In close proximity to those delay times, the probability remains high where the resulting inference curve for each exit node approximately follows a Gaussian close to the delay values $\tau_{\mathrm{node}}$, as shown in Fig. 10b. For time series where $\tau$ lays inbetween two $\tau_{\mathrm{node}}$, the exit nodes corresponding to $\tau_{\mathrm{node}}$ above and below show finite probabilities. The finer the time delay resolution in training, the better the inference results will become, up to a certain point where the network can no longer distinguish between adjacent $\tau_{\mathrm{node}}$.
@@
Figure 10: Classification probabilities for the delay time. Panel (a): Colored symbols (color intensity increasing with $\tau_{\rm node}$): The probability of inferring a delay time at an exit node corresponding to $\tau_{\mathrm{node}}$. Blue: Data corresponds to Eqs. (1) and  (2) for $k_{a}=k_{b}>0$. Red: Data corresponds to $\tanh$ feedback force (3) with $A>0$. The neural network is trained on the linear model (2) with parameter range $k_{a}\tau_{\mathrm{B}}^{2}\in[2.5,15]$, where each input trajectory is prepared as discussed in Appendix A.4. We then infer time delay for both models, linear (2)(blue) with same range of $k_{a}$ and the $\tanh$ model (3)(red) with parameter range $A\tau_{\mathrm{B}}^{2}\in[4,16]$. Average over inference results for ten different trajectories (prepared as discussed in Appendix A.4) for each parameter combination. Each data point is an average over several different values of $k_{a,b}$ and $A$ respectively. Error bars indicate one standard error of the mean. Vertical dotted lines indicate $\tau_{\rm node}$. Panel (b): Same data (but log-scaled $y$-axis) as in panel (a) represented by scaling and shifting the data to reflect the deviation from the output node $\tau_{\rm node}$. Grey Gaussian curve centered around zero as well as horizontal grey line (indicating probability one) act as guides for the eye.
@@ VI.2 Network trained on linear model applied to $\tanh$ feedback force data
For the red data points in Fig. 10, we use the same network (trained on the linear feedback force with $k_{a}=k_{b}>0$) to infer the delay time for trajectories from simulations of the $\tanh$ feedback force system, where we averaged the results for several different values of the parameter $A$ of the $\tanh$ feedback force [Eq. (3)] as well as ten noise realizations. We ensured that the linear feedback force system was trained in a parameter range that is similar to that of the $\tanh$ feedback force, if it was linearly approximated.
Overall, we find accurate inference results, with classification probabilities in excess of 80 percent for delay times close to those used in training. Looking closely, we observe that the peaks are shifted slightly to the left, that is, the network slightly overestimates the delay time in the test trajectories, especially for the longer delay times considered. Another observation concerns larger deviations that we find at the highest values of the delay time $\tau$. We interpret this observation as follows. For long delay times, the $\tanh$ feedback force (3) becomes constant at large displacements. On the other hand, for small amplitudes $A$, the change in the feedback force in response to the perturbation is no longer clearly distinguishable from thermal noise. In those cases, the neural network can no longer clearly distinguish between different delay times leading to an increased probability for small inferred delay (where $\tau=0$ corresponds to diffusive behavior).
In summary, these results indicate that a network trained in linear time-delayed feedback is able to generalize to nonlinear systems with the underlying $\tanh$ feedback force. For suitable parameter ranges, similar results can be achieved for the Gaussian feedback force discussed above. This may indicate that systems subject to time delay that behave linearly in some limit may allow classification via a neural network trained on linear feedback forces, without any further knowledge about the functional form of the feedback forces in the systems.
@@ VI.3 Network trained simultaneously on several nonlinear feedback forces
So far, we have trained our model on only one specific feedback force in a range of parameters. However, in practical applications, it may be desirable to have a more versatile network that can infer the delay time for a wide variety of feedback forces, where the underlying functional form cannot be specified with certainty. To demonstrate that, in principle, this is also possible, we trained a convolutional neural network on the different nonlinear equations of motion in Eqs. (12)-(14). The corresponding results are presented in Fig. 11. Each data point represents the classification probability averaged over two noise realizations, the three different functional forms of forces as well as several different combinations of the system parameters.
Figure 11: Classification probabilities for the delay time. Inference results for network trained on Eqs. (12)– (14), where each input trajectory is prepared as discussed in Appendix A.4. We then infer time delay for trajectories obtained from the same model. Panel (a): Symbols (color intensity increasing with $\tau_{\rm node}$): The probability of inferring a delay time at an exit node corresponding to $\tau_{\mathrm{node}}$. Average over inference results for ten different trajectories (prepared as discussed in Appendix A.4) for each parameter combination. Each data point is an average over several different values of system parameters respectively. Error bars indicate one standard error of the mean. Vertical dotted lines indicate $\tau_{\rm node}$. Panel (b): Same data (but log-scaled $y$-axis) as in panel (a) represented by scaling and shifting the data to reflect the deviation from the output node $\tau_{\rm node}$. Grey Gaussian curve centered around zero as well as horizontal grey line (indicating probability one) act as guides for the eye.
Close to the delay times on which the network was trained, we again encounter the activation curves resembling a Gaussian and inference probability around 90 percent. Further away from those delay times, we see an elevated activation of other output nodes, i.e. the probability of miss-classification increases between two values of $\tau_{\mathrm{node}}$. Although it is not possible to specify the exact reason for this behavior, it might be due to the systems dynamical response in some parameter regimes being less sensitive to the perturbation applied.
@@
Although a detailed study of the accuracy of the prediction and limitations of networks trained in general forces is desirable, it is beyond the scope of this study. Still, this proof-of-concept shows that in principle the neural network can be trained on several time-delayed feedback and instantaneous forces without sacrificing inference accuracy.
@@ VII Conclusions
Starting from linear time-delayed feedback, we have identified and analytically analyzed a characteristic signature of time delay in the PSD which carries over to nonlinear systems. On this basis we have developed a method to infer time delay and the specific delay time. The method requires no knowledge about the functional form of the of the underlying forces.
Since this PSD based method requires many long stochastic trajectories for accurate inference, we asked the question (in the view of experimental situations) whether it is possible to infer time delay based on a short observation of only few realizations of a stochastic system. To this end, we developed a perturbation-based approach that utilizes a characteristic response a system subject to time delay exhibits in the presence of an external perturbation. We then used convolutional neural networks to classify the resulting trajectory data to infer time delay for several linear and nonlinear stochastic systems subject to time delay. Besides the short required duration of the observation on the order of the time delay itself, this method, in principle, also allows inference in the absence of a non-equilibrium steady state or even in unstable systems.
We emphasize that in experiments, such a perturbation can readily be applied using state-of-the-art techniques [67, 68]. Specifically, for colloidal particles the perturbing force can be realized using optical tweezers [69, 6] or external fields [70].
In recent years several parameter inference techniques for dynamical systems and equation discovery (for ordinary as well as partial differential equations) from (noisy) data have been developed, methods ranging from Bayesian inference  [32] to the application of neural network-based techniques [33, 34]. However, to the best of our knowledge, only few such methods exist for systems with time delay, with highly limited applicability to stochastic systems. Our methods to infer time delay, which can be applied to linear as well as nonlinear systems or even without explicit knowledge of the underlying forces, substantially advance the toolbox of parameter inference in stochastic systems.
@@ Acknowledgements.
Appendix A Numerical procedure
@@
A.1 Numerical solution of the overdamped delay Langevin equation

### v2

@@ VI Perturbation and deep learning-based approach – numerical results
In the following, we discuss results of the perturbation-based approach to infer time delay using neural networks for the trajectory data obtained from Brownian dynamics simulations (see Appendix B.1 for details).
We will first focus on inference by means of a convolutional neural network trained on a linear time-delayed feedback force [see Eq. (2)]. With this network, we infer time delay for trajectory data obtained from numerical simulations of the linear feedback force [see Eq. (2)] as a proof of concept. Subsequently, the $\tanh$ feedback force [see Eq. (3)] is considered to demonstrate that the network trained on a linear feedback force is able to generalize. In principle, the perturbation-based approach to infer time delay using convolutional neural networks trained on linear feedback forces may also be applicable to other systems that exhibit a stationary or non-equilibrium steady state such as Gaussian time delayed feedback [Eq. (4)]; however, an extensive analysis is beyond the scope of this study.
Finally, we will discuss the inference results obtained from a convolutional neural network trained on trajectories based on Eqs. (11)– (13). We use the trained network to classify trajectory data obtained from numerical simulations of the same equations of motion. In this way, we demonstrate that it is possible to train such a neural network on several functional forms of the (nonlinear) delayed and instantaneous forces simultaneously while maintaining similar prediction accuracy as before. In practical applications, this allows for inference despite not knowing the exact dynamics beforehand.
@@ VI.1 Network trained on linear model applied to linear force data
For our numerical study we first consider the attractive linear feedback force system based on Eqs. (1) and  (2) with $k_{a}=k_{b}>0$. Here, we use a neural network that was trained on the same model to infer time delay. In Fig. 10a we show inference results represented by the probability distribution $P(\tau)$ (see Sec. V.2). We averaged over the classification probabilities for different values of the parameters $k_{a}$ and $k_{b}$ at each data point, as well as over $10$ noise realizations. Each of the color-coded curves corresponds to one of the exit nodes and represents the probability of a time series to be based on the delay time value $\tau_{\mathrm{node}}$. Markers correspond to the underlying delay times of test data, i.e. the time series we obtain from the perturbation approach and feed into the neural network (after applying the preparation procedure discussed in Appendix B.4). We classify time series based on delay times of $\tau=0\tau_{B}$ (passive system) up to $\tau=0.2\tau_{B}$.
@@
The highest classification probability is observed if a time series is based on a delay time close to one of the $\tau_{\mathrm{node}}$ the network was explicitly trained on. In those cases, we find $P(\tau)\approx 1$, that is, the network is almost certain. In close proximity to those delay times, the probability remains high where the resulting inference curve for each exit node approximately follows a Gaussian close to the delay values $\tau_{\mathrm{node}}$, as shown in Fig. 10b. For time series where $\tau$ lays inbetween two $\tau_{\mathrm{node}}$, the exit nodes corresponding to $\tau_{\mathrm{node}}$ above and below show finite probabilities. The finer the time delay resolution in training, the better the inference results will become, up to a certain point where the network can no longer distinguish between adjacent values of $\tau_{\mathrm{node}}$.
@@
Figure 10: Classification probabilities for the delay time. Panel (a): Colored symbols (color intensity increasing with $\tau_{\rm node}$): The probability of inferring a delay time at an exit node corresponding to $\tau_{\mathrm{node}}$. Blue: Data corresponds to Eqs. (1) and  (2) for $k_{a}=k_{b}>0$. Red: Data corresponds to $\tanh$ feedback force (3) with $A>0$. The neural network is trained on the linear model (2) with parameter range $k_{a}\tau_{\mathrm{B}}^{2}\in[2.5,15]$, where each input trajectory is prepared as discussed in Appendix B.4. We then infer time delay for both models, linear (2)(blue) with same range of $k_{a}$ and the $\tanh$ model (3)(red) with parameter range $A\tau_{\mathrm{B}}^{2}\in[4,16]$. Average over inference results for ten different trajectories (prepared as discussed in Appendix B.4) for each parameter combination. Each data point is an average over six different values of $k_{a,b}$ and seven different values of AA respectively. Error bars indicate one standard error of the mean. Vertical dotted lines indicate $\tau_{\rm node}$. Panel (b): Same data (but log-scaled $y$-axis) as in panel (a) represented by scaling and shifting the data to reflect the deviation from the output node $\tau_{\rm node}$. Grey Gaussian curve centered around zero as well as horizontal grey line (indicating probability one) act as guides for the eye.
@@ VI.2 Network trained on linear model applied to $\tanh$ feedback force data
For the red data points in Fig. 10, we use the same network (trained on the linear feedback force with $k_{a}=k_{b}>0$) to infer the delay time for trajectories from simulations of the $\tanh$ feedback force system, where we averaged the results for several different values of the parameter $A$ of the $\tanh$ feedback force [Eq. (3)] as well as $10$ noise realizations. We ensured that the linear feedback force system was trained in a parameter range that is similar to that of the $\tanh$ feedback force, if it was linearly approximated.
Overall, we find accurate inference results, with classification probabilities in excess of 80 percent for delay times close to those used in training. Looking closely at Fig. 10a, we observe that the peaks still approximately follow a Gaussian but are shifted slightly to the left, that is, the network slightly overestimates the delay time in the test trajectories, especially for the longer delay times considered. Another observation concerns larger deviations that we find at the highest values of the delay time $\tau$. We interpret this observation as follows. For long delay times, the $\tanh$ feedback force (3) becomes constant at large displacements. On the other hand, for small amplitudes $A$, the change in the feedback force in response to the perturbation is no longer clearly distinguishable from thermal noise. Therefore, the underlying reason for the observed deviations may be a dominance of diffusive behavior when the $\tanh$ feedback force assumes a constant value (beyond the linear regime) for large displacements in combination with a small feedback force amplitude compared to noise. In those cases, the neural network can no longer clearly distinguish between different delay times leading to an increased probability (see Fig. 10b) for inferring short delay (where $\tau=0$ corresponds to diffusive behavior).
In summary, these results indicate that a network trained in linear time-delayed feedback is able to generalize to nonlinear systems with the underlying $\tanh$ feedback force. For suitable parameter ranges, similar results can be achieved for the Gaussian feedback force discussed above. This indicates that systems subject to time delay that behave linearly in some limit may allow classification via a neural network trained on linear feedback forces, without any further knowledge about the functional form of the feedback forces in the systems.
@@ VI.3 Network trained simultaneously on several nonlinear feedback forces
So far, we have trained our model on only one specific feedback force in a range of parameters. However, in practical applications, it may be desirable to have a more versatile network that can infer the delay time for a wide variety of feedback forces, where the underlying functional form cannot be specified with certainty. To demonstrate that, in principle, this is also possible, we trained a convolutional neural network on the different nonlinear equations of motion in Eqs. (11)-(13). The corresponding results are presented in Fig. 11. Each data point represents the classification probability averaged over $10$ noise realizations, three different functional forms of forces as well as 27 combinations of the system parameters that are different from those used in training.
Figure 11: Classification probabilities for the delay time. Inference results for network trained on Eqs. (11)– (13), where each input trajectory is prepared as discussed in Appendix B.4. We then infer time delay for trajectories obtained from the same model. Panel (a): Symbols (color intensity increasing with $\tau_{\rm node}$): The probability of inferring a delay time at an exit node corresponding to $\tau_{\mathrm{node}}$. Average over inference results for ten different trajectories (prepared as discussed in Appendix B.4) for each parameter combination. Each data point is an average over 27 different combinations of the system parameters. Error bars indicate one standard error of the mean. Vertical dotted lines indicate $\tau_{\rm node}$. Panel (b): Same data (but log-scaled $y$-axis) as in panel (a) represented by scaling and shifting the data to reflect the deviation from the output node $\tau_{\rm node}$. Grey Gaussian curve centered around zero as well as horizontal grey line (indicating probability one) act as guides for the eye.
From Fig. 11a we observe good inference accuracy with classification probability around 90 percent close to the delay times on which the network was trained. Furthermore, we again observe activation curves that resemble a Gaussian (see Fig. 11b). Further away from those delay times, we see an elevated activation of other output nodes, i.e. the probability of miss-classification increases between two values of $\tau_{\mathrm{node}}$ but overall remains small (below 10 percent). Although it is not possible to specify the exact reason for this behavior, it might be due to the systems dynamical response in some parameter regimes being less sensitive to the perturbation applied.
@@
This proof-of-concept shows that in principle the neural network can be trained on several time-delayed feedback and instantaneous forces without sacrificing inference accuracy. Finally, in principle this method can also be used on systems with colored noise as demonstrated in Appendix D, Fig. D6.
@@ VII Conclusions
Starting from linear time-delayed feedback, we have identified and analytically analyzed a characteristic signature of time delay in the PSD, which we have shown to carry over to nonlinear systems. On this basis we have developed a method to infer time delay and the specific delay time. The method requires no knowledge about the functional form of the of the underlying forces.
Since this PSD-based method requires many long stochastic trajectories for accurate inference, we asked the question (in the view of experimental situations) whether it is possible to infer time delay based on a short observation of only few realizations of a stochastic system. To this end, we developed a perturbation-based approach that utilizes a characteristic response a system subject to time delay exhibits in the presence of an external perturbation. We then used convolutional neural networks to classify the resulting trajectory data in order to infer time delay for several linear and nonlinear stochastic systems subject to time delay. This method only requires a short duration of the observation on the order of the time delay itself. For this reason, we expect that this method, in principle, also allows inference in the absence of a non-equilibrium steady state or even in unstable systems, provided that underlying forces do not diverge during the observation time.
We emphasize that in experiments, such a perturbation can be realized in principle using state-of-the-art techniques [79, 80]. Specifically, for colloidal particles the perturbing force can be applied using optical tweezers [6, 81] or external fields [82].
In recent years several parameter inference techniques for dynamical systems and equation discovery (for ordinary as well as partial differential equations) from (noisy) data have been developed, methods ranging from Bayesian inference  [41] to the application of neural network-based techniques [43]. However, to the best of our knowledge, only few such methods exist for systems with time delay, with highly limited applicability to stochastic systems. Our methods to infer time delay, which can be applied to linear as well as nonlinear systems or even without explicit knowledge of the underlying forces, substantially advance the toolbox of parameter inference in stochastic systems.
@@
In this study, we focused on white noise systems for our analytical and numerical considerations but also demonstrated the applicability of our inference methods to systems with exponentially correlated noise. However, in real-world systems, more complex types of noise may have to be taken into consideration (e.g., different types of colored noise, multiplicative noise, spatially inhomogeneous noise). For those systems, it remains an open challenge to develop strategies to infer time delay by extending and generalizing the methods discussed here.
Furthermore, from a more fundamental perspective, it would be interesting to investigate whether one could exploit fluctuations-dissipation-like relations [85, 86, 87] to learn about the delay. We note that for the present equations of motion, there is no conventional fluctuation-dissipation relation between noise correlation and memory kernel since the systems are subject to white noise (i.e., zero correlation time), whereas the ‘memory kernel’ corresponding to discrete delay is given by a delta distribution around the delay time, $\tau$ [33, 35]. Still, it would be interesting to explore whether other types of fluctuation relations exist (e.g., when comparing positional moments [88]), and if they could provide us with information about the time delay in a stochastic system.
@@ Acknowledgements.
Appendix A Analytical details
A.1 Gaussian time-delayed feedback - analytical details
A.1.1 Expansion of the nonlinear feedback force in the limit of small displacements
For small displacements, that is, small $a$ and $\tau$, the feedback force can be linearized [62], analogously to a small delay approximation [64], resulting in a force of the form of Eq. (2), where $k_{a}=k_{b}=a/b^{2}$. Corresponding numerical results are presented in Sec. IV.3 of the main text.
A.1.2 Expansion of the nonlinear feedback force in the constant velocity state
We can exploit the marginally stable constant velocity state [62] this kind of feedback force can produce in the deterministic limit. We rewrite the equation of motion in terms of position $x(t)$ relative to the position of a particle moving at constant velocity $v_{\infty}$. We can then linearly approximate the feedback force, assuming that the deviation is small. This leaves us with an approximate equation of motion that describes the stochastic dynamics in the ‘constant velocity state’ [62]:
$\gamma\dot{x}(t)\approx\gamma v_{\infty}+K\left[x(t)-x(t-\tau)-v_{\infty}\tau\right]+\xi(t),$ (18)
where we define
$\tilde{x}_{v_{\infty}}(\omega)=\frac{(\gamma+K\tau)v_{\infty}\delta(\omega)+\tilde{\xi}}{\gamma i\omega-K+Ke^{-i\omega\tau}}.$ (21)
In our simulations, we cannot resolve the zero-frequency contribution, which we therefore neglect in the following. Under this assumption, the Fourier transform of the position would be the same as for the linear feedback force with $k_{a}=k_{b}>0$, which was discussed in the main text (see Sec, III.2). We note that persistent motion cannot be considered a stationary state but rather a non-equilibrium steady state. The variance of the position in such systems grows over time, even if the mean particle position goes to zero for long observation times or many noise realizations. For this reason, the differencing method has to be applied in order to obtain the correct numerical results. Corresponding numerical results are presented in Sec. IV.3 of the main text.
A.2 Definition of the power spectral density via correlation functions
To put our definition of the PSD [Eq. (5)] into a broader context, we note that the power spectral density is often defined via the Fourier transform of the corresponding autocorrelation function [94] under the assumption of stationarity of the underlying stochastic process. We recall that for one trajectory, the power spectral density is defined as
$\tilde{x}_{\mathcal{T}}(f)$ $=\int_{-\infty}^{\infty}x_{\mathcal{T}}(t)e^{-i2\pi ft}dt$ (23)
$=\int_{t_{0}-\mathcal{T}/2}^{t_{0}+\mathcal{T}/2}x(t)e^{-i2\pi ft}dt\ .$ (24)
$\hphantom{=}\times x_{\mathcal{T}}(t)dt\bigg]e^{-i2\pi ft^{\prime}}dt^{\prime}$ (25)
$=\int_{-\infty}^{\infty}R_{xx}(t^{\prime})e^{-i2\pi ft^{\prime}}dt^{\prime}\ ,$ (26)
with $R_{xx}$ being the autocorrelation function of the position $x(t)$. For, e.g., a confined Brownian particle, the mean and variance of the position are constant after some time. Under this condition, the power spectral density can be computed via the Fourier transform of the corresponding positional autocorrelation function as well as via the modulus of the Fourier transform of the position. We fully rely on the latter since the autocorrelation is only of limited use in analytical computations for the time delay systems considered here. Furthermore, the Fourier transform of the position of a linear system can be easily computed both analytically and numerically.
A.3 Linear time-delayed feedback - small frequency behavior
where we defined
$\tau_{\pm}\equiv\gamma\frac{1\pm\sqrt{1-k_{a}/k_{b}}}{k_{a}}\ .$ (28)
From the approximation (Eq. 27) we can see that the $\mathrm{PSD}$ scales as $1/\omega^{2}$. For $k_{a,b}>0$, this behavior can be interpreted as a reduced ability of the feedback trap to confine compared to the instantaneous harmonic trap, resulting in a corresponding deviation from the black curve in Fig. 1 at small $\omega$ due to diffusive behavior in the long time limit. Analogously, for $k_{a,b}<0$ the frequency dependency is connected to enhanced diffusion due to repulsive time-delayed feedback
A.4 Linear time-delayed feedback - marginally stable system
Appendix B Numerical procedure
@@
B.1 Numerical solution of the overdamped delay Langevin equation